OpenStudy (firejay5):

How do I get y = 12x in standard form

OpenStudy (campbell_st):

subtract y from both sides

OpenStudy (campbell_st):

what do you get

OpenStudy (firejay5):

12X - Y

OpenStudy (campbell_st):

don't forget = 0 so its 12x - y = 0 its an equation so you need the equal's sign

OpenStudy (firejay5):

is that it

OpenStudy (anonymous):

Arguably, it was in (a) standard form to start with....

OpenStudy (firejay5):

no it wasn't it was written in slope intercept form

OpenStudy (anonymous):

So why is slope intercept not a standard form?

OpenStudy (anonymous):

@estudier, indeed, there seems to be some inconsistency in mathematics between 'general forms' and 'standard forms'

OpenStudy (firejay5):

Slope intercept form is y = mx + b Standard form is AX + BY = C

OpenStudy (anonymous):

Each type of function has its own template for what it calls 'standard' and what it calls 'general.' For linear equations, the standard is Ax+By=C

OpenStudy (anonymous):

Says who?

OpenStudy (firejay5):

Says nobody

OpenStudy (anonymous):

You could try and argue that there is a "canonical" expression for a line, I don't think there is one, really.

OpenStudy (anonymous):

There are mathematical conventions that people agree on to be able to be consistent when they talk to one another. No different from having X for horizontal and Y for vertical. It's arbitrary, but we have to agree on something so we can talk to each other. There is a 'general form' for quadratics: Ax^2 +Bx +Cy^2 +Dy + Exy + F = 0 and many 'standard forms' for quadratics depending on if the graph is a parabola, circle, ellipse, etc.